A long straight wire carries a current of 1.0 A.


(a) At what distance from the wire is the magnitude of the magnetic field


equal to 2.5 x 10^-5 T?


(b)What is the magnitude of the magnetic force that this wire will exert on a


3.0 nC particle moving at 4.2 x 10^4 m/s perpendicular to the field at that


distance?

1
Expert's answer
2022-04-11T12:04:22-0400

(a)

"B=\frac{\mu_0I}{2\pi r}"

"r=\frac{\mu_0I}{2\pi B}=\frac{4\pi*10^{-7}*1.0}{2\pi*2.5*10^{-5}}=0.008\;\rm m"

(b)

"F=qvB\\
=3.0*10^{-9}*4.2*10^4*2.5*10^{-5}=3.2*10^{-9}\:\rm N"